Solve for x:
(x-4)/2-1/5=(7x+1)20

Answer by Nate(3500)

Your question:Solve for x:
(x-4)/2-1/5=(7x+1)20
-----------------------------------------------------------------------------
My answer:
(x-4)/2 -(1/5)=(7x+1)20----original equation
(x-4)/2 -(1/5)=140x+20----distribute
(x-4)/2 =140x+20.2----add 1/5
x-4=280x+40.4----multiply the equation by 2
-4=279x+40.4----get the x's all on one side
-44.4=279x----minus 40.4
-444/2790=x----get x by itself by dividing 239
-444/2790 or -74/465 is your answer!